OpenStudy (anonymous):

Factor the expression 6x^4 - 9x^3 + 15x^2 + 24x using greatest common factor

OpenStudy (jhannybean):

\[6=\color{red}{3}\cdot 2\]\[9 = \color{red}{3\cdot 3}\]\[15 = 5\cdot \color{red}{3}\]\[24 = 2\cdot 2\cdot 2 \cdot \color{red}{3}\] \(3\) is the GFC out of all the constants upon prime factoring all the terms. \(x\) is the variable that is common throughout all terms and has the smallest power. \[=3x(2x^3-3x^2+5x+8)\]